# Victor Shea-Jay Huang1,211footnotemark: 1, Le Zhuo1,211footnotemark: 1, Yi Xin2,3, Zhaokai Wang2,4, Peng Gao2, Hongsheng Li122footnotemark: 2 1CUHK, MMLab  2Shanghai AI Laboratory  3NJU  4SJTU

### v1

@@ 5.4 Diffusion Really Learned Features
5.5 Coarse-to-fine Generation in DiT
Based on our TIDE and the results shown in Fig. 5, we observed in the interpretable latent space that, from the perspective of sparse feature activation, abrupt changes in both cosine similarity and MSE metrics are concentrated at the timestep when contours are formed, indicating that this moment marks a critical transition where the latent space begins aligning the original image manifold with the target manifold. This phenomenon not only supports the hypothesis of a “coarse-to-fine” generation process but also suggests that the entire denoising process may involve a transition between two sampling modes. A similar perspective is discussed in  [11], where it highlights a potential connection between diffusion models and autoregressive models, suggesting that diffusion models essentially perform an approximate form of autoregression in the frequency domain. Specifically, the iterative denoising process in diffusion models can be interpreted as a low-pass filtering operation, where noise gradually overwhelms high-frequency signals while preserving low-frequency structures, progressively recovering image details through spectral decomposition. This frequency-domain perspective further explains the “coarse-to-fine” behavior observed during image generation in diffusion models, aligning closely with our findings from sparse feature activation, and provides additional insights into the generative logic of diffusion models.
6 Potential Application
@@
Based on the analysis in the previous section, we choose to modify the activation layer of the penultimate DiT, which has a significant impact on the output. As shown in Fig. 5, for each timestep, there always exists a directed vector between the latent feature and the target feature. At this point, the significance of Self-Attention Embedding is also reflected. By modifying the TIDE latent feature, we not only achieve better interpretability but also avoid out-of-distribution (OOD) errors caused by directly operating on the activation layer. Additionally, we can add the classifier-free guidance (CFG) Renormalization [31], as shown in the formula below, to ensure the stability of the norm.
$\tilde{F_{\theta}}(x_{t},c_{T},c_{I},t)=\frac{\hat{F}_{\theta}(x_{t},c_{T},c_{I},t)\left\|F_{\theta}(x_{t},c_{T},c_{I},t)\right\|}{\left\|\hat{F}_{\theta}(x_{t},c_{T},c_{I},t)\right\|}$ (9)
@@
TIDE-based image editing methods, while not yet matching the performance of certain highly optimized traditional approaches — such as GANs or domain-specific pre-trained models — in scenarios requiring exceptionally high-quality edits, offer distinct advantages that address key limitations of conventional techniques. Traditional methods, often reliant on prompt-based operations, can produce high-quality results but struggle with limited interpretability, insufficient control over polysemantics, and susceptibility to unpredictable outputs. In contrast, TIDE leverages sparse feature extraction and modification, enabling precise control, stability, and transparency in the editing process while providing detailed insights into the model’s internal mechanisms. These strengths make TIDE particularly valuable in applications where precision, reliability, and interpretability are critical, such as medical image modification or ensuring accountability throughout the generation process. Furthermore, TIDE extends its utility beyond image editing by guiding model optimization, reducing bias, and fostering the development of innovative generative techniques, establishing a foundation for more trustworthy technologies.
7 Limitations
Current SAE-based image editing methods still face challenges in achieving high-quality edits in complex scenarios. These limitations may stem from insufficient feature representation capacity due to the sparsity constraints of the model. Future work will focus on optimizing the design of sparsity constraints and exploring dynamic sparse activation mechanisms to improve editing quality while maintaining interpretability and stability.
8 Conclusion
This work presents a systematic exploration of interpretability in diffusion models, with a focus on DiTs. Through the introduction of the TIDE framework, we explore the unique temporal dynamics of DiTs, proposing novel training strategies, sparsity scheduling, and a decoupled analysis of the generation process. Our findings explore scaling law for TIDE, demonstrate its SoTA performance compared to other SAEs, and highlight its ability to learn hierarchical features during generative pre-training, enabling efficient transfer to downstream tasks. Furthermore, comprehensive experiments validate the practical utility of TIDE-learned sparse features in applications such as image editing and style transfer. By advancing the transparency, controllability, and interpretability of DiTs, this work provides a foundation for building more trustworthy generative systems, while also offering new directions for future research in understanding diffusion-based architectures.
[…]
Supplementary Material
A Related work
A.1 Sparse Autoencoders for Interpretability
Mechanistic interpretability [13] aims to uncover how neural networks produce their outputs by identifying the features or concepts encoded within individual neurons [40]. A key challenge arises from the polysemantic nature of neurons [12], where a single neuron can respond to multiple unrelated concepts due to superposition, a phenomenon in which neural networks encode more features than they have dimensions [12, 49]. Sparse autoencoders (SAEs) [41] address this issue by learning compact and interpretable representations through a sparsity-constrained latent space, where only a small subset of neurons activates for each input.
Recent advancements demonstrate the effectiveness of SAEs in mechanistic interpretability, enabling the discovery of human-interpretable features [22, 6] and sparse feature circuits in language models [37]. SAEs have been applied to decompose activations in large language models (LLMs), revealing features related to biases and safety that can be manipulated to steer model behavior [56, 61, 8]. Similarly, in vision-language models such as CLIP [45], SAEs have been used to extract interpretable concepts from the vision encoder [9] and facilitate image editing in diffusion models [10]. Furthermore, SAEs have enabled concept naming via embeddings from the CLIP text encoder for concept bottleneck models [46] and have been employed to disable specific generative features in text-to-image diffusion models. These applications highlight the potential of SAEs to mitigate superposition and enhance interpretability across diverse model architectures.
A.2 Interpretability of Diffusion Models
Diffusion models have shown remarkable performance in text-to-image (T2I) generation by iteratively predicting Gaussian noise within latent representations. Early implementations relied on U-Net architectures [47, 44], but recent advancements have introduced the more scalable Diffusion Transformer (DiT) [43], as demonstrated in works such as [59, 62, 7, 5, 29, 33]. Improvements in sampling efficiency [52, 51, 35, 60] have further enhanced image quality while reducing computational costs. Despite these advances, interpretability studies on diffusion models remain limited.
Efforts to interpret U-Net-based models have focused on disentangling semantic directions in bottleneck layers [28, 42, 21] and analyzing cross-attention mechanisms [55]. For example, [2, 3] demonstrated how attributes could be manipulated by adjusting text inputs at specific layers, while [58] interpreted text encoders by generating intermediate image outputs. Sparse autoencoders (SAEs), widely used in language models, have seen limited adoption in vision tasks. Early applications include manipulating CLIP representations [45, 15, 10] and traditional convolutional networks [54, 19]. More recently, SAEs have been employed in vision-language models to reduce hallucinations [24] and to enhance radiology reporting [1]. In T2I diffusion models, SAEs have been used to identify meaningful features in bottleneck layers [23], distill knowledge from SDXL-Turbo blocks [53], and analyze unconditional sampling activations across timesteps [26], linking features with visual semantics.
B Training design
B.1 Training Loss
The optimization process updates the parameters of the stacked autoencoders (SAEs) $\theta$ iteratively using the following formula:
$\theta^{(t+1)}=\theta^{(t)}-\eta_{t}\cdot\text{clip}\left(\mathbf{g}^{(t)},\tau\right)$ (10)
where $\eta_{t}$ denotes the learning rate at step $t$, and $\mathbf{g}^{(t)}$ is the gradient of the training loss with respect to the model’s parameters $\theta$. The gradient $\mathbf{g}^{(t)}$ is computed as:
$\mathbf{g}^{(t)}=\frac{\partial}{\partial\theta}\left[\ell(x,\hat{x};\theta)+\lambda_{\text{reg}}\cdot R(\theta)\right]$ (11)
Here, $\ell(x,\hat{x};\theta)$ defines the training loss, and $R(\theta)$ is the regularization term. The parameter $\lambda_{\text{reg}}$ controls the strength of regularization.
The training loss $\ell(x,\hat{x};\theta)$ consists of multiple components, each of which captures a specific aspect of the model’s performance. It is expressed as:
$\ell(x,\hat{x};\theta)=\mathcal{L}_{\text{mse}}+\mathcal{L}_{\text{cos}}+\mathcal{L}_{\text{temporal}}+\mathcal{L}_{\text{spa}}$ (12)
  • Reconstruction Loss: The reconstruction loss $\mathcal{L}_{\text{mse}}$ measures the mean squared error (MSE) between the input $x$ and its reconstruction $\hat{x}$:
    $\mathcal{L}_{\text{mse}}=\lambda_{\text{mse}}\|x-\hat{x}\|_{2}^{2}$ (13)
  • Cosine Similarity Loss: The cosine similarity loss $\mathcal{L}_{\text{cos}}$ ensures directional alignment between $x$ and $\hat{x}$ by minimizing the cosine distance:
    $\mathcal{L}_{\text{cos}}=\lambda_{\text{cos}}\left(1-\frac{\mathbf{x}\cdot\mathbf{\hat{x}}}{\|\mathbf{x}\|\|\mathbf{\hat{x}}\|}\right)$ (14)
  • Temporal Consistency Loss: The temporal consistency loss $\mathcal{L}_{\text{temporal}}$ enforces smoothness across time steps in sequential data:
    $\mathcal{L}_{\text{temporal}}=\lambda_{\text{temporal}}\|F_{t}-F_{t+1}\|_{2}^{2}$ (15)
  • Sparsity Loss: The sparsity loss $\mathcal{L}_{\text{spa}}$ promotes sparse representations in the latent space. It is implemented using the Huber loss:
    $\mathcal{L}_{\text{spa}}=\lambda_{\text{spa}}\begin{cases}\frac{1}{2}(S-S_{\text{target}})^{2}&\text{if }|S-S_{\text{target}}|\leq\delta\\ \delta|S-S_{\text{target}}|-\frac{1}{2}\delta^{2}&\text{otherwise}\end{cases}$ (16)
B.2 Regularization
To prevent overfitting and ensure well-posed optimization, the loss function includes a regularization term $R(\theta)$, defined as:
$R(\theta)=\|W_{\text{encoder}}\|_{2}^{2}+\|W_{\text{decoder}}\|_{2}^{2}$ (17)
where $W_{\text{encoder}}$ and $W_{\text{decoder}}$ denote the weights of the encoder and decoder, respectively. The $L_{2}$ norm is applied to penalize large weights, ensuring model stability and generalizability.
B.3 Gradient Clipping
To stabilize the optimization process and mitigate the risk of gradient explosion, gradient clipping is employed. The clipping function $\text{clip}(\mathbf{g},\tau)$ is defined as:
$\text{clip}(\mathbf{g},\tau)=\begin{cases}\mathbf{g},&\|\mathbf{g}\|\leq\tau,\\ \frac{\tau}{\|\mathbf{g}\|}\cdot\mathbf{g},&\|\mathbf{g}\|>\tau\end{cases}$ (18)
where $\|\mathbf{g}\|$ represents the norm of the gradient vector. When the gradient norm exceeds the threshold $\tau$, the gradient is scaled down to have a norm equal to $\tau$. This ensures that updates remain controlled, improving the stability of the convergence.
C Potential Application
Based on the analysis in the previous section, we propose a novel strategy to modify the activation layer of the penultimate DiT (Diffusion Transformer), which significantly influences the output. As illustrated in Fig.5 in the main text, for every timestep, a directed vector exists between the latent feature and the target feature. This relationship highlights the importance of the Self-Attention Embedding in achieving meaningful feature alignment. By operating on the TIDE latent feature — rather than directly on the activation layer — we not only enhance the interpretability of the process but also mitigate the risk of out-of-distribution (OOD) errors. To ensure numerical stability during these operations, we introduce classifier-free guidance (CFG) Renormalization [31], as expressed in the formula below:
$\tilde{F_{\theta}}(x_{t},c_{T},c_{I},t)=\frac{\hat{F}_{\theta}(x_{t},c_{T},c_{I},t)\cdot\left\|F_{\theta}(x_{t},c_{T},c_{I},t)\right\|}{\left\|\hat{F}_{\theta}(x_{t},c_{T},c_{I},t)\right\|},$ (19)
where:
  • $\tilde{F_{\theta}}$ represents the renormalized TIDE latent feature.
  • $\hat{F_{\theta}}$ is the modified latent feature after applying feature editing operations.
  • $F_{\theta}$ denotes the original latent feature before modification.
  • $x_{t}$ is the input at timestep $t$.
  • $c_{T}$ and $c_{I}$ represent the text and image conditions, respectively.
This renormalization ensures that the norm of the modified latent feature $\hat{F_{\theta}}$ is scaled to match the norm of the original feature $F_{\theta}$, maintaining numerical stability and preventing extreme deviations during editing. By normalizing the modified feature, we ensure that the edits remain consistent with the model’s expected latent space dynamics.
C.1 Modification Strategies
The interpretable latent space introduced by TIDE enables various modification strategies. These strategies are designed to offer fine-grained control over edits while maintaining stability:
  • Erasing or Inverting Top-K Features: By selectively erasing or inverting the top-K features of the original latent feature while preserving the overall norm, we can alter the direction of these features. This approach allows us to remove fine-grained details, as shown in the first row of Fig. 6, where intricate rococo-style elements are erased, leaving behind only the basic background and structural attributes of a house. This operation offers a baseline for more sophisticated modifications.
  • Scaling Activation Indices: By scaling the top-K activation indices of the target latent feature within the original latent feature, we can achieve smooth transitions and gradual feature changes. For example, in the second row of Fig. 6, this technique is used to progressively add aging attributes to the image of a young boy, resulting in a controlled transformation.
  • Global Replacement of Latent Features: A more aggressive editing strategy involves directly replacing the original latent feature with the target latent feature at a specific timestep. This method can produce dramatic global modifications, as illustrated in the third row of Fig. 6.
These diverse strategies showcase the flexibility of TIDE in image editing and its ability to achieve fine control over both local and global transformations.
C.2 Advantages of TIDE-Based Methods
Although TIDE-based image editing methods may not yet rival the performance of highly optimized traditional approaches — such as GANs or domain-specific pre-trained models — in tasks requiring extremely high-quality image edits, they offer several unique advantages over conventional techniques:
  • Enhanced Interpretability: Traditional methods, often reliant on prompt-based operations, struggle to fully interpret the multi-faceted semantics of individual features. Outputs are frequently influenced by unanticipated factors, making these methods less transparent. TIDE, by contrast, provides a clear understanding of the internal mechanisms of the generative model, empowering users to make informed modifications.
  • Precise Control: Through sparse feature extraction and targeted modification, TIDE enables precise control over editing outcomes. This ensures that users can achieve their desired results without compromising model stability or generating unintended artifacts.
  • Stability and Transparency: TIDE’s reliance on well-defined latent space operations ensures consistent and predictable outputs, making it a reliable tool for scenarios where accountability and trustworthiness are critical.
  • Broader Applicability: Beyond image editing, TIDE’s interpretability-focused framework has potential applications in areas such as model optimization, bias reduction, and the development of innovative generative techniques. For example, TIDE-based methods can guide the refinement of generative models or help identify and mitigate biases in generated content.
C.3 Application Scenarios
The combination of interpretability, precision, and stability makes TIDE particularly valuable in scenarios requiring high levels of control and trustworthiness:
  • Medical Image Editing: In medical imaging, where precision and accountability are paramount, TIDE’s interpretability ensures that modifications are both accurate and transparent, reducing the risk of errors in critical applications such as diagnostics or surgical planning.
  • Creative Industries: TIDE’s ability to edit fine details and achieve gradual transformations makes it a versatile tool for tasks such as artistic rendering, style transfer, or content creation.
  • Model Optimization and Bias Reduction: By offering insights into the internal workings of generative models, TIDE can guide efforts to optimize model performance and address issues such as bias or fairness in AI-generated content.

### v2

@@ Introduction
  • 3D-Level: TIDE captures features like “far,” “near,” and “shadow” that reflect foreground-background relationships. These features show highly specific activation patterns, and images with similar activations consistently share such 3D characteristics—highlighting the model’s ability to encode depth-related concepts useful for tasks like depth estimation.
  • Class-Level: TIDE learns class-specific features and their subcomponents. For example, "Rococo" activates features like "house," "window," and "door," while "eagle" activates "bird," "fly," and "beak." These activations are localized, reflecting a clear hierarchical structure in the learned representations.
  • Semantic-Level: At a finer granularity, TIDE identifies detailed semantic features and their spatial positions—e.g., pinpointing the "beak" of an eagle—demonstrating its capacity for extracting precise, localized semantic information.
As shown in Fig.1, these results confirm that TIDE effectively encodes interpretable features across multiple levels—3D, class, and semantic—within pre-trained DiTs(Chen, Yu et al. 2023). This highlights the hierarchical structure of features learned during large-scale generative pre-training, demonstrating that diffusion models inherently capture and organize multi-level concepts. Such representations support diverse downstream tasks. For example, class- and semantic-level features enable classification (Li et al. 2023; Mukhopadhyay et al. 2023) and segmentation, including zero-shot methods like DiffSeg (Tian et al. 2024). Likewise, the ability to learn 3D concepts allows these models to perform depth estimation, as shown in methods like Marigold (Ke et al. 2024) and DepthFM (Gui et al. 2024).
Table 2: The use of TIDE achieves better editing performance without compromising the baseline quality. All results are evaluated on the MS-COCO dataset using FID (↓), ASR (↑) on I2P, and AlignScore (↑) under the T2I-CompBench benchmark. The improvement in editing performance mainly stems from the noticeable increase in AlignScore, while the baseline quality remains stable as the FID only shows a marginal decrease.We highlight the best value in blue , and the second-best value in green.
[…]
Attribute Binding Object Relationship Complex
Color Shape Texture Spatial Non-Spatial
Hyper-parameters Baseline (no SAE/TIDE) 7.30 0.688 0.549 0.708 0.210 0.318 0.412
SAE (5% sparsity) 7.93 +0.63 0.602 0.479 0.652 0.148 0.254 0.352
TIDE (5%, 16d, default) 7.45 +0.15 0.647 0.514 0.679 0.185 0.271 0.392
TIDE (3% sparsity) 7.60 +0.30 0.589 0.443 0.636 0.153 0.238 0.348
TIDE (latent dim = 4d) 7.97 +0.67 0.562 0.438 0.621 0.147 0.213 0.335
TIDE (latent dim = 8d) 7.88 +0.58 0.585 0.467 0.649 0.169 0.229 0.342
TIDE (latent dim = 32d) 7.41 +0.11 0.672 0.505 0.696 0.192 0.262 0.407
DiT Layers Baseline (no SAE/TIDE) 7.30 0.688 0.549 0.708 0.210 0.318 0.412
SAE (Block 27) 7.34 +0.04 0.663 0.531 0.652 0.182 0.294 0.382
TIDE (Block 4, early) 10.49 +3.19 0.495 0.337 0.475 0.102 0.152 0.141
TIDE (Block 10, mid) 8.92 +1.62 0.607 0.441 0.554 0.133 0.256 0.244
TIDE (Block 27, penultimate) 7.64 +0.34 0.672 0.537 0.667 0.195 0.301 0.352
TIDE (Block 28, final) 7.34 +0.04 0.672 0.529 0.657 0.193 0.309 0.349
Generalization SDXL (Podell et al. 2023) 8.24 0.00 0.638 0.544 0.528 0.212 0.318 0.417
SDXL + SAE 8.44 +0.20 0.583 0.504 0.451 0.167 0.295 0.348
SDXL + TIDE 8.38 +0.14 0.607 0.528 0.495 0.181 0.309 0.373
FLUX-dev (Labs 2024) 10.09 0.00 0.612 0.583 0.480 0.208 0.287 0.401
FLUX-dev + SAE 10.18 +0.09 0.489 0.441 0.432 0.178 0.216 0.314
FLUX-dev + TIDE 10.13 +0.04 0.549 0.505 0.453 0.192 0.247 0.356
Safety SDv1.4 (Rombach et al. 2022b) 17.80 16.40 0.00 0.310 0.336 0.347 0.344 0.350 0.339
ESD (Gandikota et al. 2024) 2.84 18.08 +1.68 0.315 0.34 0.351 0.347 0.355 0.343
SA (Heng and Soh 2024) 2.78 24.96 +8.56 0.319 0.344 0.355 0.354 0.359 0.348
RECE (Gong et al. 2024) 0.80 17.02 +0.62 0.321 0.337 0.348 0.345 0.351 0.34
SAFREE (Yoon et al. 2024) 1.48 19.88 +3.48 0.316 0.341 0.342 0.348 0.356 0.344
CS (Kim and Ghadiyaram 2025) 0.42 18.59 +2.19 0.327 0.345 0.351 0.368 0.369 0.357
TIDE (ours) 0.64 19.72 +3.32 0.317 0.345 0.357 0.352 0.366 0.33
Potential Application
@@
Based on the analysis in the previous section, we choose to modify the activation layer of the penultimate DiT, which has a significant impact on the output. For each timestep, there always exists a directed vector between the latent feature and the target feature. At this point, the significance of Self-Attention Embedding is also reflected. By modifying the TIDE latent feature, we not only achieve better interpretability but also avoid out-of-distribution (OOD) errors caused by directly operating on the activation layer. Additionally, we can add the classifier-free guidance (CFG) Renormalization (Lin, Liu et al. 2024), to ensure the stability of the norm.
@@
Ablation Study
To understand the effectiveness of our proposed TIDE module, we conduct comprehensive ablation experiments across four axes: (1) hyper-parameter configurations, (2) DiT layer positioning, (3) generalization across diffusion backbones and (4) safe image editing.
Evaluation Metrics. We evaluate performance using three complementary metrics:
  • Fréchet Inception Distance (FID) (Heusel et al. 2017), using the FID-30K protocol on MS-COCO (Lin et al. 2014), to measure image generation quality.
  • Compositional Alignment, via T2I-CompBench (Huang et al. 2023), using AlignScore to evaluate semantic consistency with compositional prompts.
  • Safety, using the I2P benchmark (Schramowski et al. 2023), reporting Attack Success Rate (ASR) to assess robustness against unsafe prompt injection.
Hyper-parameter Analysis
We investigate the effect of key hyper-parameters in TIDE, including sparsity ratio, latent dimensionality, and token sampling. As shown in Table 2, although all TIDE variants exhibit slightly higher FID than the baseline, our default configuration (5% sparsity, token sampling, 16d latent) leads to the smallest increase in FID (+0.15) and least drop in AlignScore across multiple dimensions, indicating minimal performance degradation.
Lowering the sparsity to 3% further increases FID and reduces alignment, while disabling token sampling causes overfitting, reflected in a drop in both FID and semantic alignment. Increasing the latent dimension improves performance: 32d yields the highest Complex-score (0.407), though below the baseline. These results highlight a trade-off between sparsity and expressiveness, and suggest that moderate sparsity and sufficient latent capacity minimize performance impact.
Impact of DiT Layer Selection
We examine the impact of injecting TIDE at different layers of the DiT backbone. Table 2 shows that deeper layers yield better trade-offs: applying TIDE at the penultimate block (Block 27) results in the lowest FID increase (+0.34) and relatively strong alignment (e.g., 0.537 in Texture), making it the best-performing insertion point among all layers tested.
In contrast, injecting TIDE at earlier layers (e.g., 4) leads to significant performance degradation, with much higher FID (+3.19) and lower AlignScore. This confirms that deeper layers encode richer semantic features and are more suitable for sparse interpretability without harming image quality.
Generalization Across Diffusion Backbones
To assess the model-agnostic nature of TIDE, we apply it to two additional diffusion backbones: SDXL and Flux-dev. From Table 2, we observe that while FID slightly increases in all variants, TIDE introduces the smallest changes: FID increases only from 8.24 to 8.38 on SDXL, and from 10.09 to 10.13 on Flux-dev. Meanwhile, alignment performance remains close to the baseline or shows modest improvement.
These findings suggest that TIDE generalizes well across architectures, maintaining stable generation quality with low overhead. While TIDE does not improve raw generation metrics, its primary benefit lies in adding interpretability with minimal performance penalty.
Safety Evaluation
We further evaluate TIDE in the context of safe image editing using the I2P benchmark, which involves editing NSFW images to remove inappropriate content. As shown in Table 2, TIDE achieves a low Attack Success Rate (ASR) of 0.64%, competitive with state-of-the-art defenses like CS (0.42%) and SAFREE (1.48%). Additionally, TIDE maintains generation quality (FID = 19.72), outperforming many baselines in safety–quality trade-offs.
Unlike the previous experiments, this setting involves actual image editing, where TIDE must not only preserve quality but also effectively sanitize unsafe content. Our results confirm that TIDE achieves strong safety guarantees with relatively minor degradation in visual fidelity.
These ablations demonstrate that:
  • TIDE introduces minimal performance degradation, with the smallest increases in FID and smallest drops in AlignScore across hyper-parameters, layers, and backbones.
  • Moderate sparsity and 16–32d latent space offer the best trade-off between interpretability and performance.
  • Deeper DiT layers (especially Block 27) are the optimal points for interpretability injection.
  • TIDE generalizes well across diffusion backbones, confirming its robustness and portability.
  • In safe editing, TIDE filters NSFW content while preserving quality and defense.
Conclusion
TIDE introduces a sparse, temporally-aware interpretability framework for Diffusion Transformers, enabling transparent analysis of their internal representations. By uncovering multi-level semantic features, TIDE supports practical applications like image editing and demonstrates strong generalization across models and tasks. While current performance in complex editing scenarios remains limited by sparsity constraints, future work will explore more flexible sparse mechanisms to improve expressiveness. Overall, TIDE provides a solid foundation for building more trustworthy, controllable, and interpretable generative systems.